OpenStudy (anonymous):

-5x+y-4z=60 2x+4y+3z=-12 6x-3y-2z=-52 solving a 3 variable linear system

OpenStudy (andriod09):

wow. all you have to do is add them i believe.

OpenStudy (anonymous):

well my math teacher said to use elimination using two different sets of the 3 so i did but i got z as a decimal and he said all the variables were whole numbers so idk what to do

OpenStudy (andriod09):

add 2 of them, so you cab get the better of the 2, then you substitute the vaiable in the last equation.

OpenStudy (anonymous):

okay ill try, thanks

OpenStudy (anonymous):

@kirstenande this was along question!!! i got::: x=-8 y=4 z=-4

OpenStudy (andriod09):

greek, it should be y=5 not 4.

OpenStudy (anonymous):

@andriod09 substitute all of my values back into one of the 3 eqns.... LHS=RHS

OpenStudy (anonymous):

@kirstenande U getting it???

OpenStudy (andriod09):

ohhh fail.

OpenStudy (anonymous):

@andriod09 :P

OpenStudy (anonymous):

i just re did it and when i got it down to the two equations with two variables i got 22x+19z=-252 -9x-14z=128

OpenStudy (anonymous):

i just re did it and when i got it down to the two equations with two variables i got 22x+19z=-252 -9x-14z=128

OpenStudy (anonymous):

 first matrix